5. You have come across some liquid nitrogen (temperature -196ºC). You are excited to test the gas  laws you have learned in chemistry class. You blow up a balloon to a volume of 1 L at room  temperature (25ºC) and place it in the liquid nitrogen. What will the volume of the balloon be  after placing it in the liquid nitrogen?  


1
Expert's answer
2021-01-28T05:48:10-0500

V1\V2 = T1\T2

So here V1 = 1L , V2 = ? , T1 = 25 +273 = 298k , T2 = -196 +273 = 77 k

So V2 = V1 × ( T2\T1)

= 1L × ( 77\298)

= 1 L × 0.258 = 0.258 L ans.

So volume of ballon after placing liq nitrogen is 0.258L .